OpenStudy (anonymous):

Solve D = 1/4(e – 3f) for e. You must show all work.

OpenStudy (australopithecus):

\[D = \frac{1}{4(e-3f)}\] \[D(e-3f) = \frac{1}{4(e-3f)} *(e-3f)\]

OpenStudy (australopithecus):

you should be able to solve this now I assume?

OpenStudy (anonymous):

its not 1/4(e-3f the e-3f is separate from 1/4

OpenStudy (anonymous):

hellllllllppp

OpenStudy (australopithecus):

oh\[D = \frac{1}{4}(e-f)\] \[4*D= \frac{1}{4}(e-f)*4\]

OpenStudy (australopithecus):

can you solve it now?

OpenStudy (australopithecus):

sorry I'm messing around with other stuff so my responses may be delayed

OpenStudy (anonymous):

4d = e-3t 4d+3f=e ?????

OpenStudy (australopithecus):

looks good

OpenStudy (anonymous):

my test is timed :(

OpenStudy (anonymous):

can you tell me the final answer please ?

OpenStudy (australopithecus):

I replaced the variables but you are right

OpenStudy (australopithecus):

I replaced the variables because wolfram alpha is kind of sketchy with what letters you use for variables

OpenStudy (anonymous):

thankssss

OpenStudy (anonymous):

\[D\text{ =}\frac{1}{4(e-3f)}\]Cross multiply and expand the LHS.\[4 D e-12 D f = 1\]\[4 D e=1+12 D f\]\[e=\frac{1+12 D f}{4 D}\]